A restaurant charges $7 for a regular meal, and $5 for a child's meal. Both prices include tax.
a. A family has $35 to spend at the restaurant. Write an inequality to represent the maximum number of regular and child's meals they can buy.
b. Does the family have enough money to buy two regular meals and three child's meals? Explain.
Thank you in advance!

Answer by stanbon(75887)

A restaurant charges $7 for a regular meal, and $5 for a child's meal. Both prices include tax.
a. A family has $35 to spend at the restaurant. Write an inequality to represent the maximum number of regular and child's meals they can buy.
Let # of regular meals be r
Let # of child meals be c
--------
Value INEQUALITY: 7r + 5c <= 35
-----------------------------------------
b. Does the family have enough money to buy two regular meals and three child's meals? Explain
2r + 3c = 2*7 + 3*5 = 14+15 = $29
======================================
Cheers,
Stan H.